# Electron-Phonon interaction and lattice thermal conductivity from metals to 2D Dirac crystals: a review

### v1

@@ 2.2 Transition Rates, Carrier Lifetimes, and Electron-Phonon Thermal Conductivity
Starting from the e–ph Hamiltonian of the previous section, transition probabilities are obtained with Fermi’s Golden Rule[3, 4]. Because electronic dispersions, screening and carrier statistics differ sharply between metals and semiconductors, the resulting lifetimes and their impact on heat flow take two distinct forms.
In metals, electrons dominate heat transport, and their momentum-relaxation time $\tau_{e}$ is limited by e–ph scattering,
@@
where $f_{\mathbf{k}}$ and $n_{\mathbf{q}\nu}$ are the Fermi–Dirac and Bose–Einstein occupations, respectively. Within the Wiedemann–Franz picture the electronic thermal conductivity scales as
@@
with $v_{F}$ the Fermi velocity and $C_{e}=\pi^{2}k_{\mathrm{B}}^{2}TN(E_{F})/3$ the Sommerfeld specific heat[14].
@@
$\begin{split}\bigl{(}\tau^{\mathrm{e\text{-}ph}}_{\nu\mathbf{q}}\bigr{)}^{-1}=&\frac{2\pi}{\hbar}\!\sum_{mn\mathbf{k}}\bigl{|}g^{\nu}_{mn}(\mathbf{k},\mathbf{q})\bigr{|}^{2}\cdot\\ &\Big{[}f_{n\mathbf{k}}\bigl{(}1-f_{m,\mathbf{k+q}}\bigr{)}\delta\!\bigl{(}\varepsilon_{m,\mathbf{k+q}}-\varepsilon_{n\mathbf{k}}-\hbar\omega_{\nu\mathbf{q}}\bigr{)}+\bigl{(}1-f_{n\mathbf{k}}\bigr{)}f_{m,\mathbf{k+q}}\delta\!\bigl{(}\varepsilon_{n\mathbf{k}}-\varepsilon_{m,\mathbf{k+q}}-\hbar\omega_{\nu\mathbf{q}}\bigr{)}\Big{]}.\end{split}$ (10)
Inserting $\tau_{\mathbf{q}\nu}$ into the Peierls–Boltzmann expression gives[43]
@@
Although both Eqs. (8)–(10) are governed by the same underlying e–ph Hamiltonian, and share the same matrix element $g_{mn}^{\nu}(\mathbf{k},\mathbf{q})$, the relevant lifetimes and transport mechanisms depend strongly on the dominant carrier type, the density of states, and the screening properties of the material. In the next sections we will study different materials, their properties and how their thermal conductivity is affected by e-ph interaction.
@@ 3 Lattice Thermal Conductivity in Metals: The Dominant Role of Electron–Phonon Interaction
where $\kappa_{\mathrm{e}}$ and $\kappa_{\mathrm{ph}}$ (W m-1 K-1) are the electronic and lattice contributions, respectively. The additive decomposition $\kappa_{\rm tot}=\kappa_{e}+\kappa_{\rm ph}$ holds when e-ph drag is negligible. With drag, cross-coefficients appear in the coupled BTE and the simple sum is an approximation. First-principles calculations for 18 elemental and intermetallic metals show that once e-ph scattering is included, $\kappa_{\mathrm{ph}}$ may supply 1–40 % of $\kappa_{\mathrm{tot}}$ at 300 K, contradicting the common assumption that phonons are negligible in metals[9]. This is shown in Fig. 1(a).
@@
with mode label $\lambda=(\mathbf{q},\nu)$, $c_{v,\lambda}=k_{\mathrm{B}}x_{\lambda}^{2}e^{x_{\lambda}}/(e^{x_{\lambda}}-1)^{2}$, $x_{\lambda}=\hbar\omega_{\lambda}/k_{\mathrm{B}}T$, and $v^{\alpha}_{\lambda}=\partial\omega_{\lambda}/\partial q_{\alpha}$. Matthiessen’s rule gives
@@ 3.3 Klemens–Williams analytic limit
$\kappa_{\mathrm{ph}}^{\mathrm{(ph\text{-}ph+e\text{-}ph)}}=\kappa_{\mathrm{ph}}^{\mathrm{(ph\text{-}ph)}}\Bigl{[}1-\frac{\omega_{i}}{\omega_{D}}\ln\!\bigl{(}\tfrac{\omega_{D}}{\omega_{i}}+1\bigr{)}\Bigr{]},\qquad\omega_{i}=\omega_{D}\,\frac{v_{g}}{v_{F}}\,\frac{\pi n_{e}}{3B\gamma_{G}^{2}}\,\frac{D_{\mathrm{e\text{-}ph}}^{2}}{k_{\mathrm{B}}T\,E_{F}},$ (19)
@@ 3.4 Electron transport under e–ph control
with $n_{\lambda}$ the Bose occupation of phonon mode $\lambda$. Larger $\lvert g\rvert^{2}$, a denser electronic DOS near $E_{\mathrm{F}}$, or a higher thermal phonon population all shorten $\tau_{i\mathbf{k}}$, thereby reducing both $\sigma$ and $\kappa_{\mathrm{e}}$[44, 45, 46]. Ziman’s Bloch–Grüneisen analysis[44] and Grimvall’s self-energy formalism[45] remain standard references and analytical benchmarks for low-temperature power laws and the separation of momentum- and energy-relaxation times. Allen’s treatment of the Eliashberg spectral function established quantitative rules for resistivity slopes and Lorenz-number renormalisation in strongly coupled metals[46, 47].
Density-functional perturbation theory combined with Wannier interpolation marked the first-principles advance and delivers state-resolved $g^{\lambda}_{j,\mathbf{k+q};\,i,\mathbf{k}}$ on ultradense grids, enabling parameter-free predictions of $\sigma(T)$ and $\kappa_{\mathrm{e}}(T)$ for transition metals[48, 9]. These calculations indicate that e–ph scattering alone can reproduce the measured resistivity of Cu, Ag and Au in good agreement with experiment once vertex corrections are included[49].
Furthermore, numerical solvers and codes such as BoltzTraP2[50] and EPW[51] implement the linearised Boltzmann equation with onboard e–ph matrix elements, while ShengBTE extensions have begun to treat the coupled electron and phonon problems on equal footing[52]. Together these analytical and computational tools provide a seamless bridge from the long-established Bloch–Grüneisen picture to modern ab initio, mode-resolved transport in metals.
@@ 3.6 Dimensionality, screening, and electron–electron corrections
Confinement effects. Femtosecond pump–probe measurements on Au and Ag nanoparticles show that electron cooling times decrease markedly with size—from $\sim 850$ fs at diameters $\sim 25$–30 nm to $\sim 500$ fs at $\sim 3$ nm—reflecting enhanced surface scattering and modified e–ph coupling in confined geometries[53, 54, 55]. Because the dominant heat-carrying phonons in metals already have short mean-free paths, additional boundary scattering in thin films, nanowires, and grain-refined materials further suppresses the lattice contribution $\kappa_{\mathrm{ph}}$[56, 57]. This can be seen in Fig. 1(b).
@@
where the screening wave vector $k_{s}$ may be taken as the Thomas–Fermi value $k_{\mathrm{TF}}$ or, more generally, as the static/dynamic Lindhard function $k_{s}(q,\omega)$[14, 27]. Stronger (“hard”) screening (large $k_{s}$) reduces the range of long-range fields and weakens Fröhlich-like couplings; weaker (“soft”) screening has the opposite effect[44, 48].
Electron–electron corrections. For transport coefficients, the e–e contribution is sensitive to how screening is treated. Partial-wave (phase-shift) solutions of the e–e scattering problem with a consistently screened potential show that first-Born estimates overstate the e–e thermal resistivity by roughly a factor of two; using the same $k_{s}$ in both treatments restores agreement and avoids the larger ($\sim\!5\times$) discrepancies quoted in older literature[58, 59]. Accurate accounting of e–e scattering is therefore essential when deconvolving $\kappa_{\mathrm{e}}$ and $\kappa_{\mathrm{ph}}$ in nanoscale metals and in ultrafast pump–probe analyses[55, 56].
@@
Transition metals 5–30 3–18 10–40% Large $N(E_{F})$, high $\omega_{\lambda}$,
@@
*Ranges are indicative values in W m-1 K-1 at $T\!\approx\!300$ K. $\kappa_{\mathrm{ph}}^{\text{(ph-ph)}}$: only phonon–phonon scattering; $\kappa_{\mathrm{ph}}^{\text{(ph-ph+e-ph)}}$: phonon–phonon plus electron–phonon scattering. EPC $\equiv$ electron–phonon coupling; $N(E_{F})$ $\equiv$ electronic density of states at the Fermi level; $\omega_{\lambda}$ $\equiv$ phonon frequency.
The table compares typical room-temperature lattice thermal conductivities for three metal families, first assuming ph–ph scattering alone, $\kappa_{\mathrm{ph}}^{(\mathrm{ph\text{-}ph})}$, and then after e–ph damping is included, $\kappa_{\mathrm{ph}}^{(\mathrm{ph\text{-}ph+e\text{-}ph})}$. In noble metals such as Cu, Ag, and Au a low electronic density of states together with strong screening reduce the lattice channel only modestly, from $\sim$3–10 to $\sim$2–6 W m-1 K-1, so phonons carry $10\,\%$ of the total heat flow. Transition metals (Ni, Co, Pt) possess a large $N(E_{F})$, stiffer phonons, and strong deformation-potential couplings; e–ph scattering therefore cuts the lattice conductivity from $\sim$5–30 to $\sim$3–18 W m-1 K-1, leaving phonons responsible for 10–40 % of overall conduction. Ordered intermetallics such as CuAu or Cu3Au behave similarly to their parent elements but can experience additional damping because optical branches couple strongly to electrons in their multi-atom bases. Hence the combined magnitude of $N(E_{F})$, phonon frequency, and the e–ph matrix element dictates how severely the lattice channel is suppressed, ranging from almost negligible in noble metals to quantitatively significant in $d$-band and intermetallic systems.
@@ 4.1 Ab initio foundation
We use the density functional theory (DFT) to obtain the electronic eigenenergies $\epsilon_{n\mathbf{k}}$ and wave-functions $\psi_{n\mathbf{k}}$, followed by density-functional perturbation theory (DFPT) for phonon frequencies $\omega_{\nu\mathbf{q}}$ and eigendisplacements $e_{\nu\mathbf{q}}$. Within the DFT/DFPT formalism[28], the linear, self-consistent change in the Kohn–Sham potential associated with a phonon mode $(\nu,\mathbf{q})$ can be written as a mass–weighted sum of Cartesian derivatives:
@@
where $e^{\nu\mathbf{q}}_{\kappa\alpha}$ are the phonon eigenvectors and $M_{\kappa}$ the ionic masses. The corresponding electron–phonon (e–ph) matrix element is[48]
$g_{mn}^{\nu}(\mathbf{k},\mathbf{q})=\sqrt{\frac{\hbar}{2\omega_{\nu\mathbf{q}}}}\;\big{\langle}\psi_{m,\mathbf{k+q}}\big{|}\,\Delta_{\nu\mathbf{q}}V\,\big{|}\psi_{n,\mathbf{k}}\big{\rangle},$ (26)
@@
In polar crystals, recovering the correct $q\!\to\!0$ behaviour requires adding the long-range Fröhlich contribution analytically to the short-range DFPT part[48, 61]. Direct evaluation of Eq. (26) on ultra-dense $(\mathbf{k},\mathbf{q})$ grids is prohibitively expensive. Modern workflows therefore interpolate the same matrix elements from coarse DFPT meshes to dense grids using maximally localised Wannier functions (e.g. EPW)[51] or atomic-orbital interpolation[60]. This is shown in Fig. 2. Given $g_{mn}^{\nu}$, the electron-limited phonon lifetime follows from Fermi’s golden rule[48]:
$\bigl{(}\tau^{\mathrm{e\text{-}ph}}_{\nu\mathbf{q}}\bigr{)}^{-1}=\frac{2\pi}{\hbar}\sum_{mn\mathbf{k}}\bigl{|}g^{\nu}_{mn}(\mathbf{k},\mathbf{q})\bigr{|}^{2}\bigl{[}f_{n\mathbf{k}}-f_{m,\mathbf{k+q}}\bigr{]}\,\delta\!\bigl{(}\epsilon_{n\mathbf{k}}-\epsilon_{m,\mathbf{k+q}}+\hbar\omega_{\nu\mathbf{q}}\bigr{)},$ (27)
@@
with $c_{v,\nu\mathbf{q}}$ the modal heat capacity and $\mathbf{v}_{\nu\mathbf{q}}=\partial\omega_{\nu\mathbf{q}}/\partial\mathbf{q}$ the group velocity[48]. All branch-resolved trends below stem from inserting the ab initio lifetimes into this expression.

### v2

@@ 2.2 Transition Rates, Carrier Lifetimes, and Electron-Phonon Thermal Conductivity
Starting from the e–ph Hamiltonian of the previous section, transition probabilities are obtained with Fermi’s Golden Rule[taylor2002quantum, jishi2013feynman]. Because electronic dispersions, screening and carrier statistics differ sharply between metals and semiconductors, the resulting lifetimes and their impact on heat flow take two distinct forms.
In metals, electrons dominate heat transport, and their momentum‐relaxation time $\tau_{e}$ is limited by e–ph scattering,
@@
where $f_{\mathbf{k}}$ and $n_{\mathbf{q}\nu}$ are the Fermi–Dirac and Bose–Einstein occupations, respectively, and band indices are implicitly summed in the spectral delta factors. Within the Wiedemann–Franz picture, the electronic thermal conductivity scales as
@@
with $v_{F}$ the Fermi velocity and $C_{e}=\pi^{2}k_{\mathrm{B}}^{2}TN(E_{\mathrm{F}})/3$ the Sommerfeld specific heat[AshcroftMermin].
@@
$\begin{split}\bigl(\tau^{\mathrm{e\text{-}ph}}_{\nu\mathbf{q}}\bigr)^{-1}=&\frac{2\pi}{\hbar}\!\sum_{mn\mathbf{k}}\bigl|g^{\nu}_{mn}(\mathbf{k},\mathbf{q})\bigr|^{2}\cdot\\ &\Big[f_{n\mathbf{k}}\bigl(1-f_{m,\mathbf{k+q}}\bigr)\delta\!\bigl(\varepsilon_{m,\mathbf{k+q}}-\varepsilon_{n\mathbf{k}}-\hbar\omega_{\mathbf{q}\nu}\bigr)+\bigl(1-f_{n\mathbf{k}}\bigr)f_{m,\mathbf{k+q}}\delta\!\bigl(\varepsilon_{n\mathbf{k}}-\varepsilon_{m,\mathbf{k+q}}-\hbar\omega_{\mathbf{q}\nu}\bigr)\Big].\end{split}$ (11)
Inserting $\tau_{\mathbf{q}\nu}$ into the Peierls–Boltzmann expression gives[Ponce2020_RPP]
@@
Although both Eqs. (9)–(11) are governed by the same underlying e–ph Hamiltonian, and share the same matrix element $g_{mn}^{\nu}(\mathbf{k},\mathbf{q})$, the relevant lifetimes and transport mechanisms depend strongly on the dominant carrier type, the density of states, and the screening properties of the material. In the next sections we will study different materials, their properties, and how their thermal conductivity is affected by e-ph interaction.
@@ 3 Lattice Thermal Conductivity in Metals: The Dominant Role of Electron–Phonon Interaction
where $\kappa_{\mathrm{e}}$ and $\kappa_{\mathrm{ph}}$ (W m-1 K-1) are the electronic and lattice contributions, respectively. The additive decomposition $\kappa_{\rm tot}=\kappa_{e}+\kappa_{\rm ph}$ holds when e-ph drag is negligible. With drag, cross-coefficients appear in the coupled BTE and the simple sum is an approximation. First-principles calculations for 18 elemental and intermetallic metals show that once e-ph scattering is included, $\kappa_{\mathrm{ph}}$ may supply 1–40 % of $\kappa_{\mathrm{tot}}$ at 300 K, contradicting the common assumption that phonons are negligible in metals[tong2019comprehensive]. This is shown in Fig. 1(a).
@@
with mode label $\lambda\equiv(\mathbf{q},\nu)$, $c_{v,\lambda}=k_{\mathrm{B}}x_{\lambda}^{2}e^{x_{\lambda}}/(e^{x_{\lambda}}-1)^{2}$, $x_{\lambda}=\hbar\omega_{\lambda}/k_{\mathrm{B}}T$, and $v^{\alpha}_{\lambda}=\partial\omega_{\lambda}/\partial q_{\alpha}$. Matthiessen’s rule gives
@@ 3.1 Phonon–phonon scattering
While Eq. (16) describes cubic (three-phonon) processes, quartic anharmonicity generates four-phonon channels that can appreciably renormalize phonon lifetimes and reduce $\kappa_{\mathrm{ph}}$ in a number of materials. First-principles formalisms based on fourth-order interatomic force constants (IFCs) now enable direct calculation of these rates and their inclusion in the phonon BTE on equal footing with three-phonon processes [FengRuan2016_PRB, FengRuan2017_PRB]. Open-source implementations (e.g., FourPhonon within ShengBTE) make such calculations practical [HanEtAl2022_CPC_FourPhonon], and can be coupled to modern BTE solvers [CarreteEtAl2017_CPC_almaBTE, Romano2021_OpenBTE].
In practice, the total (mode-resolved) transport lifetime is
$\big(\tau_{\lambda}^{\mathrm{tot}}\big)^{-1}=\big(\tau_{\lambda}^{\mathrm{ph\text{--}ph(3)}}\big)^{-1}+\big(\tau_{\lambda}^{\mathrm{ph\text{--}ph(4)}}\big)^{-1}+\big(\tau_{\lambda}^{\mathrm{e\text{--}ph}}\big)^{-1}+\cdots,$ (17)
where $\tau_{\lambda}^{\mathrm{ph\text{--}ph(4)}}$ is assembled from IFC4 and energy/momentum-conserving combinations $\lambda\!\leftrightarrow\!\lambda_{1}\pm\lambda_{2}\pm\lambda_{3}$ [FengRuan2016_PRB, FengRuan2017_PRB, HanEtAl2022_CPC_FourPhonon].
When do four-phonon processes matter? Four-phonon scattering is well documented to be essential in several classes: First in ultrahigh-$\kappa$ materials with exceptionally large lattice thermal conductivity and stiff lattices such as diamond and graphene, where it reduces $\kappa$ by $\sim$30–60% over 300–1000 K, correcting three-phonon overestimates [FengRuan2017_PRB, HanRuan2023_PRB_Graphene]. Second, in strongly anharmonic systems, quartic terms and non-perturbative renormalization, e.g., stochastic self-consistent harmonic approximation (SSCHA), are often required to obtain stable dispersions and quantitative thermal transport [MonacelliEtAl2021_JPCM_SSCHA]. See Ref. [LindsayEtAl2018_MTP_Survey] for a broader survey of first-principles phonon transport and the evolution from three-phonon to four-phonon treatments.
However, for metals at $T\!\sim\!300$ K, phonon lifetimes are primarily limited by e-ph damping, and the dominant phonon mean free paths are typically below a few–tens of nanometers, so boundary/size effects are relevant (see, e.g., [tong2019comprehensive]). Within this metallic regime, four-phonon scattering is often a secondary correction to the e-ph limited lifetimes at room temperature, but it can become non-negligible in stiffer intermetallics, at elevated temperatures, or when optical branches play a larger role. When warranted, its inclusion is straightforward via Eq. (17) by adding $\tau_{\lambda}^{\mathrm{ph\text{--}ph(4)}}$ to the total scattering rate.  [^3] In Section 3 (metals, $T\!\sim\!300$ K) we model lattice thermal transport using ph-ph scattering limited to cubic (three-phonon) processes together with explicit e-ph damping; four-phonon (quartic) scattering is not included in the presented results.
@@ 3.3 Klemens–Williams analytic limit
$\kappa_{\mathrm{ph}}^{\mathrm{(ph\text{-}ph+e\text{-}ph)}}=\kappa_{\mathrm{ph}}^{\mathrm{(ph\text{-}ph)}}\Bigl[1-\frac{\omega_{i}}{\omega_{D}}\ln\!\bigl(\tfrac{\omega_{D}}{\omega_{i}}+1\bigr)\Bigr],\qquad\omega_{i}=\omega_{D}\,\frac{v_{g}}{v_{F}}\,\frac{\pi n_{e}}{3B\gamma_{G}^{2}}\,\frac{D_{\mathrm{e\text{-}ph}}^{2}}{k_{\mathrm{B}}T\,E_{\mathrm{F}}},$ (21)
@@ 3.4 Electron transport under e–ph control
with $n_{\lambda}$ the Bose occupation of phonon mode $\lambda$. Larger $\lvert g\rvert^{2}$, a denser electronic DOS near $E_{\mathrm{F}}$, or a higher thermal phonon population all shorten $\tau_{i\mathbf{k}}$, thereby reducing both $\sigma$ and $\kappa_{\mathrm{e}}$[Ziman1960, Grimvall1981, Allen1975_PRB]. Ziman’s Bloch–Grüneisen analysis[Ziman1960] and Grimvall’s self-energy formalism[Grimvall1981] remain standard references and analytical benchmarks for low-temperature power laws and the separation of momentum- and energy-relaxation times. Allen’s treatment of the Eliashberg spectral function established quantitative rules for resistivity slopes and Lorenz-number renormalisation in strongly coupled metals[Allen1975_PRB, Allen1987_RMP].
Density-functional perturbation theory combined with Wannier interpolation marked the first-principles advance and delivers state-resolved $g^{\lambda}_{j,\mathbf{k+q};\,i,\mathbf{k}}$ on ultradense grids, enabling parameter-free predictions of $\sigma(T)$ and $\kappa_{\mathrm{e}}(T)$ for transition metals[Giustino2017_RMP, tong2019comprehensive]. These calculations indicate that e–ph scattering alone can reproduce the measured resistivity of Cu, Ag and Au in good agreement with experiment once vertex corrections are included[Mustafa2016_PRB].
Furthermore, numerical solvers and codes such as BoltzTraP2[Madsen2018_CPC] and EPW[Ponce2016_CPC] implement the linearised Boltzmann equation with onboard e–ph matrix elements, while ShengBTE extensions have begun to treat the coupled electron and phonon problems on equal footing[Li2020_PRL]. Together these analytical and computational tools provide a seamless bridge from the long-established Bloch–Grüneisen picture to modern ab initio, mode-resolved transport in metals.
@@ 3.6 Dimensionality, screening, and electron–electron corrections
Confinement effects. Femtosecond pump–probe measurements on Au and Ag nanoparticles show that electron cooling times decrease markedly with size—from $\sim 850$ fs at diameters $\sim 25$–30 nm to $\sim 500$ fs at $\sim 3$ nm—reflecting enhanced surface scattering and modified e–ph coupling in confined geometries[Sun1994_PRB, DelFatti2000_ChemPhys, Hartland2011_ChemRev]. Because the dominant heat‐carrying phonons in metals already have short mean-free paths, additional boundary scattering in thin films, nanowires, and grain-refined materials further suppresses the lattice contribution $\kappa_{\mathrm{ph}}$[Cahill2003_JAP, Majumdar1993_JHT]. This can be seen in Fig. 1(b).
@@
where the screening wave vector $k_{s}$ may be taken as the Thomas–Fermi value $k_{\mathrm{TF}}$ or, more generally, as the static/dynamic Lindhard function $k_{s}(q,\omega)$[AshcroftMermin, kazemian2023dynamic]. Stronger (“hard”) screening (large $k_{s}$) reduces the range of long-range fields and weakens Fröhlich-like couplings; weaker (“soft”) screening has the opposite effect[Ziman1960, Giustino2017_RMP].
Electron–electron corrections. For transport coefficients, the e–e contribution is sensitive to how screening is treated. Partial-wave (phase-shift) solutions of the e–e scattering problem with a consistently screened potential show that first-Born estimates overstate the e–e thermal resistivity by roughly a factor of two; using the same $k_{s}$ in both treatments restores agreement and avoids the larger ($\sim\!5\times$) discrepancies quoted in older literature[QuinnFerrell1958, Gurzhi1968_Usp]. Accurate accounting of e–e scattering is therefore essential when deconvolving $\kappa_{\mathrm{e}}$ and $\kappa_{\mathrm{ph}}$ in nanoscale metals and in ultrafast pump–probe analyses[Hartland2011_ChemRev, Cahill2003_JAP].
@@
Transition metals 5–30 3–18 10–40% Large $N(E_{\mathrm{F}})$, high $\omega_{\lambda}$,
@@
*Ranges are indicative values in W m-1 K-1 at $T\!\approx\!300$ K. $\kappa_{\mathrm{ph}}^{\text{(ph-ph)}}$: only phonon–phonon scattering; $\kappa_{\mathrm{ph}}^{\text{(ph-ph+e-ph)}}$: phonon–phonon plus electron–phonon scattering. EPC $\equiv$ electron–phonon coupling; $N(E_{\mathrm{F}})$ $\equiv$ electronic density of states at the Fermi level; $\omega_{\lambda}$ $\equiv$ phonon frequency.
The table compares typical room-temperature lattice thermal conductivities for three metal families, first assuming ph–ph scattering alone, $\kappa_{\mathrm{ph}}^{(\mathrm{ph\text{-}ph})}$, and then after e–ph damping is included, $\kappa_{\mathrm{ph}}^{(\mathrm{ph\text{-}ph+e\text{-}ph})}$. In noble metals such as Cu, Ag, and Au a low electronic density of states together with strong screening reduce the lattice channel only modestly, from $\sim$3–10 to $\sim$2–6 W m-1 K-1, so phonons carry $10\,\%$ of the total heat flow. Transition metals (Ni, Co, Pt) possess a large $N(E_{\mathrm{F}})$, stiffer phonons, and strong deformation-potential couplings; e–ph scattering therefore cuts the lattice conductivity from $\sim$5–30 to $\sim$3–18 W m-1 K-1, leaving phonons responsible for 10–40 % of overall conduction. Ordered intermetallics such as CuAu or Cu3Au behave similarly to their parent elements but can experience additional damping because optical branches couple strongly to electrons in their multi-atom bases. Hence the combined magnitude of $N(E_{\mathrm{F}})$, phonon frequency, and the e–ph matrix element dictates how severely the lattice channel is suppressed, ranging from almost negligible in noble metals to quantitatively significant in $d$-band and intermetallic systems.
@@ 4.1 Ab initio foundation
We use the density functional theory (DFT) to obtain the electronic eigenenergies $\epsilon_{n\mathbf{k}}$ and wave-functions $\psi_{n\mathbf{k}}$, followed by density-functional perturbation theory (DFPT) for phonon frequencies $\omega_{\mathbf{q}\nu}$ and eigendisplacements $e_{\nu\mathbf{q}}$. Within the DFT/DFPT formalism[baroni2001phonons], the linear, self-consistent change in the Kohn–Sham potential associated with a phonon mode $(\nu,\mathbf{q})$ can be written as a mass–weighted sum of Cartesian derivatives:
@@
where $e^{\nu\mathbf{q}}_{\kappa\alpha}$ are the phonon eigenvectors and $M_{\kappa}$ the ionic masses. The corresponding electron–phonon (e–ph) matrix element is[Giustino2017_RMP]
$g_{mn}^{\nu}(\mathbf{k},\mathbf{q})=\sqrt{\frac{\hbar}{2\omega_{\mathbf{q}\nu}}}\;\big\langle\psi_{m,\mathbf{k+q}}\big|\,\Delta_{\nu\mathbf{q}}V\,\big|\psi_{n,\mathbf{k}}\big\rangle,$ (28)
@@
In polar crystals, recovering the correct $q\!\to\!0$ behaviour requires adding the long-range Fröhlich contribution analytically to the short-range DFPT part[Giustino2017_RMP, Verdi2015_PRL]. Direct evaluation of Eq. (28) on ultra-dense $(\mathbf{k},\mathbf{q})$ grids is prohibitively expensive. Modern workflows therefore interpolate the same matrix elements from coarse DFPT meshes to dense grids using maximally localised Wannier functions (e.g. EPW)[Ponce2016_CPC] or atomic-orbital interpolation[Agapito2018_PRB]. This is shown in Fig. 2. Given $g_{mn}^{\nu}$, the electron-limited phonon lifetime follows from Fermi’s golden rule[Giustino2017_RMP]:
$\bigl(\tau^{\mathrm{e\text{-}ph}}_{\nu\mathbf{q}}\bigr)^{-1}=\frac{2\pi}{\hbar}\sum_{mn\mathbf{k}}\bigl|g^{\nu}_{mn}(\mathbf{k},\mathbf{q})\bigr|^{2}\bigl[f_{n\mathbf{k}}-f_{m,\mathbf{k+q}}\bigr]\,\delta\!\bigl(\epsilon_{n\mathbf{k}}-\epsilon_{m,\mathbf{k+q}}+\hbar\omega_{\mathbf{q}\nu}\bigr),$ (29)
@@
with $c_{v,\nu\mathbf{q}}$ the modal heat capacity and $\mathbf{v}_{\nu\mathbf{q}}=\partial\omega_{\mathbf{q}\nu}/\partial\mathbf{q}$ the group velocity[Giustino2017_RMP]. All branch-resolved trends below stem from inserting the ab initio lifetimes into this expression.